Calculate the position of Fermi level Ef at 200K for germanium crystal containing

10x1022 Boron atom per meter cube.


1
Expert's answer
2022-03-29T16:46:28-0400

"E_f=kT\ln\frac{N_d}{N_p}=1.46\cdot 10^{-19}~J."